OpenStudy (anonymous):

Which of the following is a solution |x-8| less than or equal to 14

OpenStudy (anonymous):

\[|x-8|\le14\]

OpenStudy (anonymous):

A. x = 6 B. x = 22 C. x = –4 D. x = 0

OpenStudy (anonymous):

those are my choices

OpenStudy (anonymous):

so, that would be giving you either that\[x-8\le14\]or, \[x-8\ge-14\] so your solution set (or the equation to find it) would be given this way:

OpenStudy (anonymous):

\[-14 \le x-8 \le 14\]

OpenStudy (anonymous):

now, add 8 to each piece\[-14 \color{red}{+8}\le x-8\color{red}{+8} \le 14\color{red}{+8}\] what do you get, can you tell me?

OpenStudy (anonymous):

-6 22

OpenStudy (anonymous):

16

OpenStudy (anonymous):

my answer wld be 22 rite

OpenStudy (anonymous):

the answer would be \[-6 \le x \le 22\]

OpenStudy (anonymous):

saying that x can be any [real] number from (and including) -6 ---> till (and including) 22.

OpenStudy (anonymous):

In interval notation this can be written as\[x~ \color{red}{\in}~[-6,22]\]

OpenStudy (anonymous):

thank you